OpenStudy (anonymous):

Sin10/cos370 - tan10

OpenStudy (shamim):

i can try to solve it

OpenStudy (anonymous):

That'd be amazing!

OpenStudy (shamim):

\[\frac{ \sin10 }{ \cos370 }-\frac{ \sin10 }{ \cos10 }

OpenStudy (shamim):

i think u know cos370=cos(360+10)=cos10

OpenStudy (shamim):

then we can write\[\frac{ \sin10 }{ \cos370 }-\frac{ 10 }{ \cos10 }\]

OpenStudy (shamim):

\[=\frac{ \sin10 }{ \cos10 }-\frac{ \sin10 }{ \cos10 }=0\]

OpenStudy (anonymous):

Oh that makes sense thank you!

OpenStudy (shamim):

welcome

OpenStudy (shamim):

medal